OpenStudy (anonymous):

238 = -7 ( 6 -5k) help please???

OpenStudy (anonymous):

erica just do exactly same thing like I told u before .. multiply -7(6-5k) first ..

OpenStudy (anonymous):

238 = -42 + 35k 238+42 =35k 280/35=k

OpenStudy (anonymous):

I see where I messed up I had a negative 42 still when it was suppose to be a positive

OpenStudy (anonymous):

when -42 going to lest side it become +42 :)

OpenStudy (anonymous):

left side*

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

good luck :)